1. How has the breakeven point in number of sales tickets (number of customer orders written) and break-even sales in dollars changed from 2003, to 2004, and to 2006? How has the margin of safety changed? What caused the changes?

2. One idea that the consultant had was to reduce prices to bring in more customers. If average prices were reduced ten percent (10%), and the number of sales tickets (unit sales) increased to 7,500, would the company's income be increased? With prices reduced, what would be the new break-even point in sales tickets and sales dollars? (answer 2 pls)Exhibit 2 Hallstead Jewelers Operating Statistics 2003 2004 2006 Sales space (square feet) Sales per square foot Sales ticketExhibit 1 Hallstead Jewelers; Income Statements for Years Ended January 31 (thousands of dollars) 2003 2004 2006 $8,583 |4,32rices reduced, what would be the new break-even point in sales tickets and sales dollars?
Answer #1

(thousand doller) 2003 2004 2006 sales 8,583 8102 10,711 less: Livanable cost cost of good sold 4326 4132 5570 commission - 4Break can even chane 1 2003 to 2004 to 2006 2004 = 500o - 4535=465 a number of ticket) 2006 = 7504-5000= 2504 (Momberor ticke2003 2004 2006 contmbletion mase in Ratio Go.55% 60. 30 44.58 - Break even point 3679-429 $3768 - 405 55947-56 in pollar Go.